the velocity of sound in a gas is 498 m/s and in air is 332 m/s . What is the ratio of the wavelength od sound waves in the gas to air ?

Grade:11

1 Answers

Arun
6 years ago
Dear Siddhant
 

Velocity= velocity of sound in air/velocity of sound in medium.

velocity= 332/498

Since frequency= c/lamda

frequency is proportional to
1/lamda



velocity=lamda2/lamda1

velocity=498/332

velocity=12/8

Velocity=4/3

the ratio is 4:3.
 
Regards
Arun (askIITians forum expert)